OpenStudy (anonymous):

what is the equation, in slope-intercept form, of the line that passes through (0,3) and has a slope of -6?

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Then solve for y

OpenStudy (triciaal):

that is point- slope format slope -intercept y = m x + b

pooja195 (pooja195):

I know, you can use the point slope formula to get slope intercept form...

OpenStudy (anonymous):

Still confused

pooja195 (pooja195):

wait u dont need that

pooja195 (pooja195):

y=mx+b <---slope intercept form m=slope (x,y) since they already gave the y value which is 3 u just plug it into the equation m=-6 y=3 y=mx+b <---plug in the given info :)

OpenStudy (anonymous):

So 3=-6x+?

OpenStudy (anonymous):

0?

pooja195 (pooja195):

no b=3... sorry >.< y=-6x+____ <---b value

OpenStudy (anonymous):

Okay thanks lol so y=-6x+3

OpenStudy (triciaal):

hate to but in don't want to confuse anyone but x = 0

OpenStudy (triciaal):

3 = -6*0 + b so b = 3 then y = -6x + 3

pooja195 (pooja195):

still the same answer...

pooja195 (pooja195):

Thanks for explaining better though :-)

OpenStudy (anonymous):

Thanks for explaining clearly pooja195

OpenStudy (anonymous):

Thanks for confusing me a little trical

OpenStudy (anonymous):

Help me with another ?

OpenStudy (triciaal):

really? I just pointed out that we are replacing the x = 0 that was given .

OpenStudy (anonymous):

Just messing with you

pooja195 (pooja195):

That wasn't nice.

pooja195 (pooja195):

Apologize.

OpenStudy (anonymous):

I was joking

OpenStudy (anonymous):

I apologizing for joking

OpenStudy (anonymous):

Apologize*

OpenStudy (anonymous):

Juan is saving money for a new car. He has already saved 200$ and plans to save $25 per week. This situation can be represented by the function f(x)=200+25x, where x is the number of weeks. How much money has Juan saved after 8 weeks?

OpenStudy (triciaal):

given x = 8 substitute in the formula given